# High-impedance resonators for strong coupling to an electron on helium

### v1

@@ III.1 Resonator design and fabrication
[…]
@@
We fabricate the resonators following the process developed in Ref. [21] using thin films of TiN deposited on Si (100) via plasma-enhanced atomic layer deposition at 270 C. We perform 150 cycles of atomic layer deposition (ALD), corresponding to an approximate film thickness of 12 nm. TiN can grow in the (200) and (111) crystallographic orientations, where the (200) orientation is associated with higher quality factors and lower microwave losses [31]. To verify the chemical composition of our devices, we perform X-ray diffraction (XRD) spectroscopy on both the TiN film and on the bare silicon substrate (Fig. 3d). The background-subtracted intensity displays a Gaussian diffraction peak at $2\Theta=42.5^{\circ}$, corresponding to the (200) orientation of TiN [36, 31, 21, 37, 32, 35], while no notable signal from TiN (111) is observed. The (200) orientation is consistent with other works where TiN is grown at elevated temperatures ($>$200C) with differing deposition techniques (see Table 1) [38, 36, 31, 35, 34], suggesting that temperature-related film strain and surface energy, and not the Si substrate orientation, are the key determining factors for the preferred film orientation [31, 38, 36, 39].
The resulting sheet inductance $L_{\square}$ is found from a four point probe resistance measurement across a narrow TiN strip, which reveals a critical temperature of $T_{c}=2.80\pm 0.05$ K and sheet resistance of $R_{\square}=361\pm 26$ $\Omega/\square$ (Fig  3d). From these measurements, we determine a sheet inductance $L_{\square}=\hbar R_{\square}/1.76\pi k_{B}T_{c}=178\pm 13$ pH/$\square$. Compared to previous TiN films fabricated with ALD on Si (111) [21], our devices have a lower $T_{c}$ which indicates a higher degree of disorder and results in higher overall sheet inductance $L_{\square}$.
@@ III.2 Resonator microwave characterization
The internal quality factors of all resonators exceed $Q_{i}\sim 10^{5}$, and the median internal quality factor of both Device A and Device B is $Q_{i}=3.9\times 10^{5}$, corresponding to a median resonator internal loss rate $\kappa_{i}/2\pi=11.7$ kHz. Since the $Q_{i}$ statistics for device A and B are similar, we conclude that the presence of the mock electrode and airbridge (Fig. 3c) does not limit $Q_{i}$ for both devices. Furthermore, it suggests that both devices have a common origin limiting $Q_{i}$, such as dielectric loss introduced during fabrication of both devices, or losses exacerbated by oxidation of TiN [41, 42].
@@ III.3 Resonator model verification
To quantitatively determine the extent to which the SCR circuit model describes the experimental resonators, we extract the capacitance matrix using finite element modelling software, and determine the inductances from the wire geometry and measured sheet inductance. Next, we diagonalize the circuit Hamiltonian and compare the predicted eigenfrequencies with the measured frequencies (see App. F and Table 2 for details). We find that taking the raw values of the capacitance matrix yields inaccurate predicted eigenfrequencies, because the distributed capacitance of the meandering wire does not perfectly reflect the circuit model. To address this, we discount the simulated capacitances for all resonators by a single parameter $\gamma<1$, which accounts for the distributed nature of the capacitance [44, 45, 21]. By comparing measured and simulated eigenfrequencies, and treating $\gamma$ as a fit parameter we find $<2$% error for all 18 resonators provided $\gamma=0.61\pm 0.01$ (Fig. 3b,c). The small error, combined with the lack of structure in the error confirms the SCR circuit model accurately represents the fabricated resonators.
@@ IV.   Discussion
Studying the interaction of electrons on helium with SCRs requires the addition of helium microchannels to shuttle electrons towards the dot, and electrodes that control the electron motional frequency via tuning the dot electrostatic potential. Such electrodes may cause a decrease in quality factor if they cannot be placed along a symmetry axis, especially for the high impedance SCR design [16], which can be mitigated by effective on-chip low-pass filters [48]. In addition, effective tuning of the dot curvature necessitates biasing the resonator center pin. To do this, the SCR tail can terminate in an electrode connected to a dc voltage source. This dc connection would act as a resistive component to the effective circuit that damps and totally suppresses the common mode. Experiments involving resonators with integrated microchannels, biasing electrodes, and coupling of electrons to the differential mode are currently underway.
@@ Appendix A Resonator Lagrangian with inductive tail
Figure 5: (a) Application of the Y-$\Delta$ transform for the circuit of Fig. 1, including the inductive tail. The transformed circuit is analytically tractable using the Lagrangian method, and includes an inductive coupling $L_{3}$. (b) Measured and simulated mode splitting for the 9 resonators. Circuit model refers to the simulation procedure outlined in App. F. Eq. (A8) is an approximation of this procedure which assumes equal inductances and capacitances.
@@ Appendix D Resonator fabrication
After ALD, the resonator geometry is defined using a combination of optical lithography and etching. We perform optical lithography with a Heidelberg MLA150 Direct Writer to define the TiN etch mask. After lithography, we etch the TiN using a chlorine-based Plasma-Therm inductively coupled plasma etch recipe.
@@ Appendix E Experimental setup
Appendix F Modeling of resonator properties
@@
$f_{0}=\frac{1}{2\pi\sqrt{L_{\mathrm{tot}}C_{\mathrm{tot}}}}=\frac{1}{4\sqrt{(L_{\ell}\ell)(C_{\ell}\ell)}},$ (F47)
where $L_{\mathrm{tot}}$, $C_{\mathrm{tot}}$ are the inductance and capacitance participating in the resonance, and $L_{\ell}$, $C_{\ell}$ are the effective inductance and capacitance per unit length. Note that $\ell$ is half the total resonator length (see Fig. 7a), $L_{\ell}$ is found from Eq. (F46) and $C_{\ell}\ell$ is the simulated capacitance $C_{\mathrm{sim}}$ found from Ansys Electronics Desktop. Therefore, in the case of a straight $\lambda/2$ resonator we have
$C_{\mathrm{tot}}=\left(\frac{2}{\pi}\right)^{2}C_{\ell}\ell=\gamma\,C_{\mathrm{sim}},$ (F48)

### v2

@@ III.1 Resonator design and fabrication
Figure 3: Experimental realization of SCRs (a) Scanning electron micrograph of a fabricated TiN (dark gray) resonator on Si substrate (light gray). The wire width $w=1.6$ $\mu$m. (b) To test how control electrodes affect the resonance frequency and $Q_{i}$, we test an additional realization (Device B), which contains a mock electrode and air bridge along the resonator symmetry axis (c) Layout of a test chip containing nine resonators of varying meander length $\ell$. The resonators are coupled to a common microwave feedline to measure $S_{21}$. (d) X-ray diffraction spectrum for a 100 nm thick TiN film grown on Si (100) only shows evidence of the TiN (200) crystallographic orientation. The intensity was corrected for Si (100) background counts. (e) dc resistance measured across a test structure shows $T_{c}=2.80\pm 0.05$ K (vertical dashed line). Using the measured sheet resistance $R_{\square}=361\pm 26$ $\Omega$ (horizontal dashed line) we find $L_{\square}=178\pm 13$ pH/$\square$. (f) Sample resonance for one of the resonators on device A with $f_{0}=5.025$ GHz. (g) Cumulative histogram of internal quality factors of 18 measured resonances for device A (blue, median $Q_{i}=3.9\times 10^{5}$) and device B (green, median $Q_{i}=3.9\times 10^{5}$). This shows $Q_{i}$ remains high even after adding a mock electrodes.
@@
We fabricate the resonators following the process developed in Ref. [21] using thin films of TiN deposited on Si (100) via plasma-enhanced atomic layer deposition at 270 C. We perform 150 cycles of atomic layer deposition (ALD), corresponding to an approximate film thickness of 12 nm. Further fabrication details can be found in Appendix D.
TiN can grow in the (200) and (111) crystallographic orientations, where the (200) orientation is associated with higher quality factors and lower microwave losses [31]. To verify the chemical composition of our devices, we perform X-ray diffraction (XRD) spectroscopy on both the TiN film and on the bare silicon substrate (Fig. 3d). The background-subtracted intensity displays a Gaussian diffraction peak at $2\Theta=42.5^{\circ}$, corresponding to the (200) orientation of TiN [36, 31, 21, 37, 32, 35], while no notable signal from TiN (111) is observed. The (200) orientation is consistent with other works where TiN is grown at elevated temperatures ($>$200C) with differing deposition techniques (see Table 1) [38, 36, 31, 35, 34], suggesting that temperature-related film strain and surface energy, and not the Si substrate orientation, are the key determining factors for the preferred film orientation [31, 38, 36, 39].
The resulting sheet inductance $L_{\square}$ is found from a four point probe resistance measurement across a narrow TiN strip, which reveals a critical temperature of $T_{c}=2.80\pm 0.05$ K and sheet resistance of $R_{\square}=361\pm 26$ $\Omega/\square$ (Fig  3e). From these measurements, we determine a sheet inductance $L_{\square}=\hbar R_{\square}/1.76\pi k_{B}T_{c}=178\pm 13$ pH/$\square$. Compared to previous TiN films fabricated with ALD on Si (111) [21], our devices have a lower $T_{c}$ which indicates a higher degree of disorder and results in higher overall sheet inductance $L_{\square}$.
@@ III.2 Resonator microwave characterization
The internal quality factors of all resonators exceed $Q_{i}\sim 10^{5}$, and the median internal quality factor of both Device A and Device B is $Q_{i}=3.9\times 10^{5}$, corresponding to a median resonator internal loss rate $\kappa_{i}/2\pi=11.7$ kHz. For more details on the power dependence of $Q_{i}$ see Appendix F. Since the $Q_{i}$ statistics for device A and B are similar, we conclude that the presence of the mock electrode and airbridge (Fig. 3c) does not limit $Q_{i}$ for both devices. Furthermore, it suggests that both devices have a common origin limiting $Q_{i}$, such as dielectric loss introduced during fabrication of both devices, or losses exacerbated by oxidation of TiN [41, 42].
@@ III.3 Resonator model verification
To quantitatively determine the extent to which the SCR circuit model describes the experimental resonators, we extract the capacitance matrix using finite element modelling software, and determine the inductances from the wire geometry and measured sheet inductance. Next, we diagonalize the circuit Hamiltonian and compare the predicted eigenfrequencies with the measured frequencies (see App. G and Table 2 for details). We find that taking the raw values of the capacitance matrix yields inaccurate predicted eigenfrequencies, because the distributed capacitance of the meandering wire does not perfectly reflect the circuit model. To address this, we discount the simulated capacitances for all resonators by a single parameter $\gamma<1$, which accounts for the distributed nature of the capacitance [44, 45, 21]. By comparing measured and simulated eigenfrequencies, and treating $\gamma$ as a fit parameter we find $<2$% error for all 18 resonators provided $\gamma=0.61\pm 0.01$ (Fig. 4b,c). The small error, combined with the lack of structure in the error confirms the SCR circuit model accurately represents the fabricated resonators. We note that the common-differential mode splitting changes sign at increased resonator lengths due to the decreasing ratio of the tail inductance to the total inductance (see Appendix A for details).
@@ IV.   Discussion
It is possible to increase the resonator impedance further by reducing the wire width $w$. From the scaling of $w$ and $\ell$ (Fig. 4d), we estimate that resonators with similar resonance frequency ($\omega_{0}^{\prime}=\omega_{0}$), with half the length ($\ell^{\prime}=\ell/2$) and an 8×8$\times$8 × reduced width ($w^{\prime}=w/8$) give a
$\frac{Z^{\prime}}{Z}=\frac{\ell^{\prime}}{\ell}\frac{w}{w^{\prime}}\frac{\omega_{0}^{\prime}}{\omega_{0}}=4$ (11)
@@
Studying the interaction of electrons on helium with SCRs requires the addition of helium microchannels to shuttle electrons towards the dot, and electrodes that control the electron motional frequency via tuning the dot electrostatic potential. Such electrodes may cause a decrease in quality factor if they cannot be placed along a symmetry axis, especially for the high impedance SCR design [16], which can be mitigated by effective on-chip low-pass filters [48]. In addition, effective tuning of the dot curvature necessitates biasing the resonator center pin. To do this, the SCR tail can terminate in an electrode connected to a dc voltage source. This dc connection would act as a resistive component to the effective circuit that damps and totally suppresses the common mode. In this case, any asymmetry in the resonator mode profile also results in a reduction of $Q_{i}$ for the differential mode. We estimate that a 1% asymmetry in the capacitances causes the differential mode $Q_{i}$ to drop below $10^{5}$. Experiments involving resonators with integrated microchannels, biasing electrodes, and coupling of electrons to the differential mode are currently underway.
@@ Appendix A Resonator Lagrangian with inductive tail
Figure 5: (a) Application of the Y-$\Delta$ transform for the circuit of Fig. 1, including the inductive tail. The transformed circuit is analytically tractable using the Lagrangian method, and includes an inductive coupling $L_{3}$. (b) Measured and simulated mode splitting for the 9 resonators. Circuit model refers to the simulation procedure outlined in App. G. Eq. (A8) is an approximation of this procedure which assumes equal inductances and capacitances. In this equation, $C$ includes any coupling capacitance to the feedline, which can be found in Table 2.
@@ Appendix C Coupling strength $g$ for the SCR design
We note that Eq. (C44) only includes coupling to $\dot{\phi}_{d}$ in the case of symmetric capacitances $C_{a}=C_{b}=C$. For $C_{a}\neq C_{b}$, $\mathcal{H}_{\mathrm{e-res}}$ also includes a coupling to the common mode. This term is proportional to $\dot{\phi}_{c}$. For the asymmetries in this work, we estimate that the common mode coupling can become an appreciable fraction of the differential mode coupling. In future devices the common mode coupling can be suppressed through symmetric feedline coupling. Finally, any single-electron decoherence resulting from this coupling can be suppressed by increasing the common-differential mode splitting with a larger tail inductance $L_{t}$.
@@ Appendix D Resonator fabrication
After ALD, the resonator geometry is defined using a combination of optical lithography and etching. We perform optical lithography with a Heidelberg MLA150 Direct Writer to define the TiN etch mask. After lithography, we etch the TiN using a chlorine-based Plasma-Therm inductively coupled plasma etch recipe. The etch recipe combines Cl2:BCl3:Ar with flow rates 30:30:10 sccm, and uses 400 W ICP power and 50 W bias power at a 5 mTorr chamber pressure. The resist mask remaining after TiN etching was removed in AZ NMP rinse at 80 C, followed by rinsing with IPA.
@@ Appendix E Experimental setup
Appendix F Probe power dependence of $Q_{i}$
We show the power dependence of all device A resonances in Fig. 7. All $Q_{i}$ are extracted below the bifurcation power. The mean photon number $\bar{n}$ was inferred from [51]
$\bar{n}=\frac{P_{\mathrm{in}}Q_{c}}{\hbar\omega_{0}^{2}}\left(\frac{Q_{i}}{Q_{i}+Q_{c}}\right)^{2},$ (F46)
which takes into account the power at the sample, $P_{\mathrm{in}}$, which is estimated from the VNA output power and the line attenuation of approx. 75 dB. Apart from the common mode of R9, all resonances only show a limited power dependence and have quality factors above $10^{5}$.
Figure 7: Extracted quality factors vs. resonator photon occupation $\bar{n}$. The quality factors are color-coded by resonator (R1-R9) and divided by mode type (common vs. differential). See Appendix. F for more details.
Appendix G Modeling of resonator properties
@@
$f_{0}=\frac{1}{2\pi\sqrt{L_{\mathrm{tot}}C_{\mathrm{tot}}}}=\frac{1}{4\sqrt{(L_{\ell}\ell)(C_{\ell}\ell)}},$ (G48)
where $L_{\mathrm{tot}}$, $C_{\mathrm{tot}}$ are the inductance and capacitance participating in the resonance, and $L_{\ell}$, $C_{\ell}$ are the effective inductance and capacitance per unit length. Note that $\ell$ is half the total resonator length (see Fig. 8a), $L_{\ell}$ is found from Eq. (G47) and $C_{\ell}\ell$ is the simulated capacitance $C_{\mathrm{sim}}$ found from Ansys Electronics Desktop. Therefore, in the case of a straight $\lambda/2$ resonator we have
$C_{\mathrm{tot}}=\left(\frac{2}{\pi}\right)^{2}C_{\ell}\ell=\gamma\,C_{\mathrm{sim}},$ (G49)